OpenStudy (anonymous):

What would be equivalent to (4 + 5i)(4 - 5i)?

OpenStudy (anonymous):

foil it

OpenStudy (anonymous):

or , you can use the formula (a+b)(a-b)=a^2-b^2

OpenStudy (anonymous):

so, (4 + 5i)(4 - 5i) = 4^2-(5i)^2 =16- 25i^2

OpenStudy (anonymous):

can you go from there?

OpenStudy (anonymous):

multiply a complex number by its complex conjugate, you get real number \[(a+bi)(a-bi)=a^2+b^2\]

OpenStudy (anonymous):

oh wow, i"ve actually never learned the foil method this is new to me lol

OpenStudy (anonymous):

don't think about \(i\) and don't think about minus signs just \(4^2+5^2\)

OpenStudy (anonymous):

hmm, that's to confusing

OpenStudy (anonymous):

really? foiling is confusing, sum of squares is pythagoras

OpenStudy (anonymous):

i would say use your knowleg about real algebra and keep in mind that i*i = -1

OpenStudy (anonymous):

Ok, thanks you guys :)